OpenStudy (anonymous):

Multiply (5√2-15√5)(4√2+12√5) I got (√27+√230)(√18+√149) √486+√4023+√4140+√34,270 Is this the final answer or can it be simplified.

OpenStudy (anonymous):

if they do not have a common radical then yes it is.

OpenStudy (mertsj):

\[(5\sqrt{2}-15\sqrt{5})(4\sqrt{2}+12\sqrt{5})=40+60\sqrt{10}-60\sqrt{10}-900=-860\]